# version 0.6

### v3

@@ 7.3 The Gradient Perturbation Controller
After exploring the relationships between different classes of controllers, we are ready to prove the main new algorithm in non-stochastic control, which gives a strong regret guarantee against adversarially chosen perturbations.
The algorithm is described in figure (3), for time varying linear dynamical systems. It is a specialization of the template introduced in Algorithm 1 from Section 1.6.2. The main idea is to learn the parametrization of a DAC policy using known online convex optimization algorithms, namely online gradient descent. We henceforth prove that this parametrization is essentially convex, allowing us to prove the main regret bound.
In algorithm (3) we denote by $\mathbf{x}_{t}({M}_{1:h})$ the hypothetical state arrived at by playing a DAC policy ${M}_{1:h}$ from the beginning of time; note that this may not be equal to the true state $\mathbf{x}_{t}$ of the dynamical system at time $t$. The counterfactual $\mathbf{u}_{t}({M}_{1:h})$ is defined similarly.
1:  Input: $h$, $\eta$, initialization $M_{1:h}^{1}$.
2:  Observe the linear system $(A_{t},B_{t})$, compute a stabilizing linear controller sequence ${K}_{t}$.
@@
7:     Construct loss $\ell_{t}({M}_{1:h})=c_{t}(\mathbf{x}_{t}({M}_{1:h}),\mathbf{u}_{t}({M}_{1:h})).$Update ${M}_{1:h}^{t+1}\leftarrow{M}_{1:h}^{t}-\eta\nabla\ell_{t}({M}_{1:h}^{t}).$
@@
The GPC algorithm comes with a very strong performance guaranteee: namely it guarantees vanishing worst case regret vs. the best DAC policy in hindsight. As per the relationship between DAC and other poicies, this implies vanishing regret vs. LDC, GLC and linear controllers. The formal statement is given as follows.
@@
    The costs $c_{t}$ are convex, bounded, and have bounded gradients w.r.t. both arguments $\mathbf{x}_{t}$ and $\mathbf{u}_{t}$.
@@
    The matrices $\{{A}_{t},{B}_{t}\}$ are stabilizable, and have bounded $\ell_{2}$ norm.
@@
$\max_{\mathbf{w}_{1:T}:\mathopen{\|}\mathbf{w}_{t}\mathclose{\|}\leq 1}\left(\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})-\min_{\pi\in\Pi^{DAC}}\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t}^{\pi},\mathbf{u}_{t}^{\pi})\right)\leq\mathcal{O}(\sqrt{T}).$
Furthermore, the time complexity of each loop of the algorithm is linear in the number of system parameters and logarithmic in T.
Proof sketch.
As per Theorem A.1 in Appendix A, the online gradient descent algorithm for convex loss functions ensures that the regret is $\mathcal{O}(\sqrt{T})$. However, to directly apply the OGD algorithm to our setting, we need to make sure that the following 3 conditions hold true:
@@
  • The loss function is time dependent because the loss depends on the current state, which further depends on the previous states. The state at each time step depends on the current variable estimates. Hence, $\mathbf{x}_{t}$ can be shown to be a function of $\left\{{M}^{i}_{1:h}\right\}_{i=1}^{t}$, where ${M}^{i}_{1:h}$ denotes the variables at step $i$. Thus, we need to ensure that optimizing the loss function is similar to optimizing a function which doesn’t have such a dependence on time.
First, we show that the loss function is convex w.r.t. to the variables ${M}_{1:h}$. This follows since the states and the controls are linear transformations of the variables.
@@
Proof.
Using $\mathbf{x}_{t}$ to denote $\mathbf{x}_{t}(M_{1:h})$ and $\mathbf{u}_{t}$ to denote $\mathbf{u}_{t}(M_{1:h})$, the loss function $\ell_{t}$ is given by
$\ell_{t}({M}_{1:h})=c_{t}(\mathbf{x}_{t},\mathbf{y}_{t}).$
@@
$\mathbf{x}_{t+1}={A}_{t}\mathbf{x}_{t}+{B}_{t}\mathbf{u}_{t}+\mathbf{w}_{t}$ $={A}_{t}\mathbf{x}_{t}+{B}_{t}\left({K}_{t}\mathbf{x}_{t}+\sum_{i=1}^{h}{M}_{i}\mathbf{w}_{t-i}\right)+\mathbf{w}_{t}$
$=\tilde{A}_{t}\mathbf{x}_{t}+\left({B}_{t}\sum_{i=1}^{h}{M}_{i}\mathbf{w}_{t-i}+\mathbf{w}_{t}\right),$
where $\tilde{A}_{t}=A_{t}+B_{t}K_{t}$. By induction, we can further simplify
$\mathbf{x}_{t+1}$ $=\sum_{i=0}^{t}\left(\prod_{j=t-i+1}^{t}\tilde{A}_{j}\right)\left(B_{t-i}\sum_{j=1}^{h}{M}_{j}\mathbf{w}_{t-i-j}+\mathbf{w}_{t-i}\right),$ (7.1)
which is a linear function of the variables. In the above computation, for convenience, we have assumed that the initial state $\mathbf{x_{0}}$ is $\mathbf{0}$.
@@
$\mathbf{u}_{t}={K}_{t}\mathbf{x}_{t}+\sum_{i=1}^{h}{M}_{i}\mathbf{w}_{t-i},$
and since $\mathbf{x}_{t}$ was shown to be linear in the variables, so is $\mathbf{u}_{t}$. Thus, we have shown that $\mathbf{x}_{t}$ and $\mathbf{u}_{t}$ are linear transformations of ${M}_{1:h}$ and hence, the loss function is convex in ${M}_{1:h}$. (Why? See Exercise 7.5.) ∎
Now, by the regret bound implied by OGD, we have
$\sum_{t=1}^{T}l_{t}(M_{1:h}^{t})-\min_{M_{1:h}}\sum_{t=1}^{T}l_{t}(M_{1:h})\leq\mathcal{O}(\sqrt{T}).$ (7.2)
Finally, the actual loss $c_{t}$ at time $t$ is not calculated on $\mathbf{x}_{t}(M_{1:h})$, but rather on the true state $\mathbf{x}_{t}$, which in turn depends on different parameters $M_{1:h}^{i}$ for various historical times $i<t$. Nevertheless, $c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$ is well approximated by $l_{t}(M^{t}_{1:h})$, as shown next.
Lemma 7.4.
For all $t$, $|l_{t}(M^{t}_{1:h})-c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})|\leq\mathcal{O}(1/\sqrt{T})$.
Proof.
First, note that to achieve $\mathcal{O}(\sqrt{T})$ regret, OGD uses a step size of $\eta=1/\sqrt{T}$. Since the gradients of the loss functions are bounded, we have
$\|M_{1:h}^{t}-M_{1:h}^{t-i}\|\leq\sum_{s=t-i+1}^{t}\|M_{1:h}^{s}-M_{1:h}^{s-1}\|\leq\mathcal{O}\left(\frac{i}{\sqrt{T}}\right).$
Now, we use the fact that the iterates of OGD $M_{1:h}^{t}$ move slowly to establish that $\mathbf{x}_{t}$ and $\mathbf{x}_{t}(M^{t}_{1:h})$ are close in value. Similar to the expansion of $\mathbf{x}_{t}(M_{1:h})$ in Equation 7.1, $\mathbf{x}_{t}$ can be written as
$\mathbf{x}_{t+1}$ $=\sum_{i=0}^{t}\left(\prod_{j=t-i+1}^{t}\tilde{A}_{j}\right)\left(B_{t-i}\sum_{j=1}^{h}{M}^{t-i}_{j}\mathbf{w}_{t-i-j}+\mathbf{w}_{t-i}\right).$
Recall that the stability of the time-varying linear policy $\{K_{t}\}$ implies that for any $i\leq j$, $\|\prod_{t=i}^{j}\tilde{A}_{t}\|\leq\kappa(1-\delta)^{j-i}$. Therefore, there exists a constant $C\geq 0$ such that
$\|\mathbf{x}_{t+1}-\mathbf{x}_{t}(M_{1:h}^{t})\|$ $\leq\sum_{i=0}^{t}\left\|\left(\prod_{j=t-i+1}^{t}\tilde{A}_{j}\right)\right\|\|B_{t-i}\|\sum_{j=1}^{h}\|{M}^{t-i}_{j}-M^{t}_{j}\|\|\mathbf{w}_{t-i-j}\|$
$\leq Ch\eta\sum_{i=0}^{t}i(1-\gamma)^{i}$
$\leq\frac{Ch\eta}{\gamma^{2}}=\mathcal{O}\left(\frac{1}{\sqrt{T}}\right).$
This display also implies that $\|\mathbf{u}_{t}-\mathbf{u}_{t}(M_{1:h}^{t})\|=\mathcal{O}\left({1}/{\sqrt{T}}\right)$, since $\mathbf{u}_{t}$ and $\mathbf{u}_{t}(M_{1:h}^{t})$ only differ in the immediate state they operate on. By definition, $l_{t}(M_{1:h}^{t})=c_{t}(\mathbf{x}_{t}(M_{1:h}^{t}),\mathbf{u}_{t}(M_{1:h}^{t}))$. Finally, we note that $c_{t}$ has bounded derivatives finishing the proof of the lemma. ∎
Now, we combine the regret guarantee in equation 7.2 and the immediately preceding lemma as
$\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})-\min_{\pi\in\Pi^{DAC}}\sum_{t=1}^{T}c_{t}(\hat{\mathbf{x}}_{t},\pi(\hat{\mathbf{x}}_{t}))$
$\leq\sum_{t=1}^{T}l_{t}(M_{1:h}^{t})-\min_{\pi\in\Pi^{DAC}}l_{t}(M_{1:h}^{\pi})+T\times\mathcal{O}\left(\frac{1}{\sqrt{T}}\right)=\mathcal{O}(\sqrt{T}),$
implying the stated claim.
@@ 7.4 Bibliographic Remarks
The bulk of the material of this chapter is from [ABH+19]. The work of [ABH+19] defined the general online nonstochastic setting in which the disturbances and the cost functions are adversarially chosen, and where the cost functions are arbitrary convex functions. Follow up work by [AHS19] achieves logarithmic pseudo-regret for strongly convex, adversarially selected losses, and well-conditioned stochastic noise. Logarithmic regret algorithms were derived for online nonstochastic control in [FS20].
Online nonstochastic control with known dynamics was further extended in several directions. Control with bandit feedback for the costs was considered in [GHH20]. More refined regret metrics applicable to changing dynamics were studied in [GHM20, MGSH21, ZWZ22, Sin22]. [LDL21] studies algorithms that balance regret minimization while adhering to constraints. An application of nonstochastic control in iterative planning in the presence of perturbations is described in [AHMS21]. Deluca [GHM20], a control library, provides benchmarked implementations of many such algorithms.
@@
The roots of online nonstochastic control can be traced to the online stochastic control setting, which we cover in Chapter 10.
@@
A precursor to nonstochastic control is a line of work, beginning with [AYBK14, CHK+18], focused on the control of stochastic systems under sequentially revealed adversarial costs functions. See [CK20, CCK22] for further developments.

### v4

@@ 7.3 The Gradient Perturbation Controller
After motivating the framework, defining regret, and exploring the relationships between different policy classes, we are ready to prove the main new algorithm in non-stochastic control.
The algorithm is described in figure (4), for time varying linear dynamical systems. It is a specialization of the template introduced in Algorithm 1 from Section 1.6.2. The main idea is to learn the parametrization of a DAC policy using known online convex optimization algorithms, namely online gradient descent. We henceforth prove that this parametrization is convex, allowing us to prove the main regret bound.
The notation we use to describe the algorithm is as follows. We denote by $\mathbf{x}_{t}({M}_{1:h})$ the hypothetical state arrived at by playing a DAC policy ${M}_{1:h}$ from the beginning of time. Note that this may not be equal to the true state $\mathbf{x}_{t}$ of the dynamical system at time $t$. The counterfactual $\mathbf{u}_{t}({M}_{1:h})$ is defined similarly.
1:  Input: $h$, $\eta$, initialization $M_{1:h}^{1}\in\mathcal{K}$.
2:  Observe the linear system $(A_{t},B_{t})$, compute a stabilizing linear controller ${K}_{t}$ as per Theorem 4.4.
@@
7:     Construct loss $\ell_{t}({M}_{1:h})=c_{t}(\mathbf{x}_{t}({M}_{1:h}),\mathbf{u}_{t}({M}_{1:h})).$Update ${M}_{1:h}^{t+1}\leftarrow\prod_{\mathcal{K}}\left[{M}_{1:h}^{t}-\eta\nabla\ell_{t}({M}_{1:h}^{t})\right].$
@@
The GPC algorithm comes with a very strong performance guaranteee: it guarantees vanishing regret vs. the best Disturbance Action Control policy belonging to our reference class in hindsight. As per the relationship between DAC and other policies, this implies vanishing regret vs. linear controllers as well. The formal statement is given as follows.
@@
    The costs $c_{t}$ are convex, bounded, and $L$-Lipschitz.
@@
    The matrices $\{{A}_{t},{B}_{t}\}$ are $\gamma$-stabilizable, and have bounded $\ell_{2}$ norm.
@@
$\max_{\mathbf{w}_{1:T}:\mathopen{\|}\mathbf{w}_{t}\mathclose{\|}\leq 1}\left(\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})-\min_{\pi\in\Pi^{D}}\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t}^{\pi},\mathbf{u}_{t}^{\pi})\right)\leq\mathcal{O}(LGD\gamma^{2}h\sqrt{T}),$
where the $O$ notation hides constants that only depend on the norm of $A_{t},B_{t}$. Furthermore, the time complexity of each loop of the algorithm is linear in the number of system parameters and logarithmic in T.
Note: we henceforth assume that the linear dynamical systems $(A_{t},B_{t})$ are a linear time-invariant system $(A,B)$ which is $\gamma$-stable, and there is no need for a stabilizing controller $K_{t}=0$. The extension to time-varying stabilizable systems is left as exercise 7.5.
Proof.
As per Theorem 6.1, the online gradient descent algorithm for convex loss functions ensures that the regret is bounded by $2GD\sqrt{T}$. However, to directly apply the OGD algorithm to our setting, we need to ensure that the following conditions hold true:
@@
  • We need to ensure that optimizing the loss function $\ell_{t}$ as a function of the parameters $M_{1:h}$ is similar to optimizing the cost $c_{t}$ as a function of the state and control.
These facts are formalize in the following two lemmas, that we state here and prove at the end of this section.
@@
Lemma 7.4.
Let $G$ denote a upper bound on the gradients of the $L$-Lipschitz losses $c_{t}$, and $D$ denote the diameter of the matrices $M_{1:h}$. Then there exists a constant $C$ that depends only on $\|A\|,\|B|\|$, such that for all $t$,
$|l_{t}(M^{t}_{1:h})-c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})|\leq\frac{CL\gamma^{2}hD}{\sqrt{T}}.$
The regret bound of GPC can be derived from these lemmas as follows. Note that $G$ is an upper bound on the gradients of $\ell_{t}$ as a function of the parameters $M_{1:h}$, while $L$ is the Lipschitz constant of $c_{t}$ as a function of the state and control.
$\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})-\min_{\pi\in\Pi^{D}}\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t}^{\pi},\mathbf{u}_{t}^{\pi})$
$\leq\sum_{t=1}^{T}l_{t}(M_{1:h}^{t})-\min_{\pi\in\Pi^{DAC}}l_{t}(M_{1:h}^{\pi})+T\times\frac{L\gamma^{2}ChD}{\sqrt{T}}$ Lemma 7.4
$\leq 2GD\sqrt{T}+{\sqrt{T}L\gamma^{2}ChD}$ Theorem 6.1
$\leq O(GDL\gamma^{2}h\sqrt{T}).$
7.3.1 Convexity of the losses
To conclude the analysis, We first show that the loss functions are convex w.r.t. to the variables ${M}_{1:h}$. This follows since the states and the controls are linear transformations of the variables.
Proof of Lemma 7.3.
For simplicity, assume that the initial state $\mathbf{x_{0}}$ is $\mathbf{0}$. Using $\mathbf{x}_{t}$ to denote $\mathbf{x}_{t}(M_{1:h})$ and $\mathbf{u}_{t}$ to denote $\mathbf{u}_{t}(M_{1:h})$, the loss function $\ell_{t}$ is given by
$\ell_{t}({M}_{1:h})=c_{t}(\mathbf{x}_{t},\mathbf{u}_{t}).$
@@ 7.3.1 Convexity of the losses
$\mathbf{x}_{t+1}={A}\mathbf{x}_{t}+{B}\mathbf{u}_{t}+\mathbf{w}_{t}$ $={A}\mathbf{x}_{t}+{B}\left(\sum_{i=1}^{h}{M}_{i}\mathbf{w}_{t-i}\right)+\mathbf{w}_{t}$
By induction, we can further simplify
$\mathbf{x}_{t+1}$ $=\sum_{i=0}^{t}{A}^{i}\left(B\sum_{j=1}^{h}{M}_{j}\mathbf{w}_{t-i-j}+\mathbf{w}_{t-i}\right),$ (7.2)
which is a linear function of the variables.
@@
$\mathbf{u}_{t}=\sum_{i=1}^{h}{M}_{i}\mathbf{w}_{t-i}.$
Thus, we have shown that $\mathbf{x}_{t}$ and $\mathbf{u}_{t}$ are linear transformations of ${M}_{1:h}$. A composition of convex and linear functions is convex (see Exercise 7.5), which concludes our Lemma. ∎
7.3.2 Loss functions with memory
The actual loss $c_{t}$ at time $t$ is not calculated on $\mathbf{x}_{t}(M_{1:h})$, but rather on the true state $\mathbf{x}_{t}$, which in turn depends on different parameters $M_{1:h}^{i}$ for various historical times $i<t$. Nevertheless, $c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$ is well approximated by $l_{t}(M^{t}_{1:h})$, as shown next.
Proof of Lemma 7.4.
We first note that by the choice of step size by the Online Gradient Descent algorithm 3, as per Theorem 6.1, we have that $\eta=\frac{D}{G\sqrt{T}}$. Thus,
$\|M_{1:h}^{t}-M_{1:h}^{t-i}\|\leq\sum_{s=t-i+1}^{t}\|M_{1:h}^{s}-M_{1:h}^{s-1}\|\leq i\eta G=\frac{iD}{\sqrt{T}}.$
We next use this fact to establish that $\mathbf{x}_{t}$ and $\mathbf{x}_{t}(M^{t}_{1:h})$ are close. Similar to the expansion of $\mathbf{x}_{t}(M_{1:h})$ in Equation 7.2, $\mathbf{x}_{t}$ can be written as
$\mathbf{x}_{t+1}$ $=\sum_{i=0}^{t}A^{i}\left(B\sum_{j=1}^{h}{M}^{t-i}_{j}\mathbf{w}_{t-i-j}+\mathbf{w}_{t-i}\right).$
Recall that the stability of the time-invariant linear dynamical system $\{A,B\}$ implies that for any $i$, $\|A^{i}\|\leq\kappa(1-\delta)^{i}$, with $\frac{\kappa}{\delta}\leq\gamma$. Therefore, there exists a constant $C\geq 0$ such that
$\|\mathbf{x}_{t}-\mathbf{x}_{t}(M_{1:h}^{t})\|$ $=\left\|\sum_{i=0}^{t}A^{i}B\sum_{j=1}^{h}{M}^{t-i}_{j}\mathbf{w}_{t-i-j}-\sum_{i=0}^{t}A^{i}B\sum_{j=1}^{h}M^{t}_{j}\mathbf{w}_{t-i-j}\right\|$
$\leq\sum_{i=0}^{t}\left\|A^{i}\right\|\|B\|\sum_{j=1}^{h}\|{M}^{t-i}_{j}-M^{t}_{j}\|\|\mathbf{w}_{t-i-j}\|$
$\leq C\cdot\kappa\sum_{i=0}^{t}(1-\delta)^{i}\cdot i\frac{hD}{\sqrt{T}}$
$=\frac{Ch\kappa D}{\sqrt{T}}\sum_{i=1}^{\infty}i(1-\delta)^{i-1}$ $\sum_{i=1}^{\infty}i(1-\delta)^{i}=\frac{1}{\delta^{2}}$
$\leq\frac{C\gamma^{2}hD}{\sqrt{T}}.$
The control at time $t$ depends only on the parameters at the current iteration and not on historical parameters, hence $\mathbf{u}_{t}=\mathbf{u}_{t}(M_{1:h}^{t})$. By definition, $l_{t}(M_{1:h}^{t})=c_{t}(\mathbf{x}_{t}(M_{1:h}^{t}),\mathbf{u}_{t}(M_{1:h}^{t}))$. Thus, we have
$l_{t}(M_{1:h}^{t})-c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$ $=c_{t}(\mathbf{x}_{t}(M_{1:h}^{t}),\mathbf{u}_{t}(M_{1:h}^{t}))-c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$
$\leq L\cdot\|\mathbf{x}_{t}(M_{1:h}^{t})-\mathbf{x}_{t}\|\leq\frac{L\gamma^{2}ChD}{\sqrt{T}}.$
@@ 7.4 Bibliographic Remarks
The bulk of the material of this chapter is from [ABH+19]. The work of [ABH+19] defined the general online nonstochastic setting in which the disturbances and the cost functions are adversarially chosen, and where the cost functions are arbitrary convex functions. Follow up work by [AHS19] achieves logarithmic pseudo-regret for strongly convex, adversarially selected losses, and well-conditioned stochastic noise. Logarithmic regret algorithms were derived for online nonstochastic control in [Sim20, FS20].
Online nonstochastic control with known dynamics was further extended in several directions. Control with bandit feedback for the costs was considered in [GHH20]. More refined regret metrics applicable to changing dynamics were studied in [GHM20, MGSH21, ZWZ22, Sin22]. More extensions suitable for changing environments were studied in [BW22].
[LDL21] studies algorithms that balance regret minimization while adhering to constraints. An application of nonstochastic control in iterative planning in the presence of perturbations is described in [AHMS21]. Deluca [GHM20], a control library, provides benchmarked implementations of many such algorithms.
@@
A precursor to online nonstochastic control is a line of work, beginning with [AYBK14, CHK+18], focused on the control of stochastic systems under sequentially revealed adversarial costs functions. See [CK20, CCK22] for further developments.